# Rapid Adoption, Hidden Risks: The Dual Impact of Large Language Model Customization

### v2

@@ Abstract
The increasing demand for customized Large Language Models (LLMs) has led to the development of solutions like GPTs. These solutions facilitate tailored LLM creation via natural language prompts without coding. However, the trustworthiness of third-party custom versions of LLMs remains an essential concern. In this paper, we propose the first instruction backdoor attacks against applications integrated with untrusted customized LLMs (e.g., GPTs). Specifically, these attacks embed the backdoor into the custom version of LLMs by designing prompts with backdoor instructions, outputting the attacker’s desired result when inputs contain the pre-defined triggers. Our attack includes 3 levels of attacks: word-level, syntax-level, and semantic-level, which adopt different types of triggers with progressive stealthiness. We stress that our attacks do not require fine-tuning or any modification to the backend LLMs, adhering strictly to GPTs development guidelines. We conduct extensive experiments on 4 prominent LLMs and 5 benchmark text classification datasets. The results show that our instruction backdoor attacks achieve the desired attack performance without compromising utility. Additionally, we propose an instruction-ignoring defense mechanism and demonstrate its partial effectiveness in mitigating such attacks. Our findings highlight the vulnerability and the potential risks of LLM customization such as GPTs.
@@ Introduction
To address this challenge, transformative solutions like GPTs [2] (and counterparts from various LLM providers [3]) have emerged, enabling users to develop personalized LLMs through natural language prompts without the need for programming skills. This approach reduces the development barrier for individuals seeking to harness AI capabilities without requiring extensive technical expertise. More importantly, these GPTs can be shared with others. The popularity of this approach is evident. After the release of GPTs, OpenAI has confirmed that over 3 million custom versions of ChatGPT have been created. [^1] https://openai.com/blog/introducing-the-gpt-store
@@
While the primary focus revolves around creating impactful GPTs, an essential concern remains on the trustworthiness [54] of third-party GPTs. Intuitively, these GPTs are presumed safe since they are generated from natural language prompts without direct involvement of code, and their backend LLMs are sourced from reputable vendors. Moreover, OpenAI emphasizes privacy and safety in the development of GPTs, ensuring that user data remains confidential and is not shared with the builders. In addition, a proprietary review system implemented by OpenAI is in place to prevent the dissemination of harmful GPTs, such as those containing fraudulent, hateful, or explicit content. Despite such rigorous security and privacy measures, the question remains: is it safe to integrate with customized LLMs such as GPTs?
@@
Our Work. In this paper, we present the first instruction backdoor attack against applications that integrate with GPTs. Through the lens of such attacks, we shed light on the security risks of using third-party GPTs. To our best knowledge, previous research on backdoor attacks, including those against LLMs [26], resolves around the training-time setting. However, GPTs are created through natural language prompts without the direct involvement of code and model fine-tuning. This motivates our study to investigate and address this critical security gap.
@@
Methodology. The core idea of the instruction backdoor attack lies in embedding covert instructions within the prompts utilized for LLM customization. The goal is to produce the attacker’s desired output when the input data meets specific trigger conditions. Our attack can be categorized into three levels, i.e., word, syntax, and semantic-level attacks. Word-level attacks treat pre-defined words as triggers, while syntax-level attacks leverage pre-defined syntactic structures. Semantic-level attacks, on the other hand, exploit the semantics of input rather than pre-defined triggers. To enhance the efficacy of semantic-level attacks, we incorporate Chain of Thought (CoT) [63] when constructing task instructions, facilitating LLMs to better execute backdoor instructions. These varied attack levels offer increasing levels of stealthiness. Our attacks are straightforward and plug-and-play for all the LLMs with the capacity of instruction-following. Furthermore, we propose a potential defense strategy involving the insertion of an ignoring instruction before the input, which partially mitigates the impact of backdoor instructions.
@@
Evaluation. We conduct extensive experiments involving four popular LLMs, namely LLaMA2 [56], Mistral [31], Mixtral [32], GPT-3.5 [15], along with five benchmark text classification datasets, including Stanford Sentiment Treebank (SST-2) [53], SMS Spam (SMS) [12], AGNews [68], DBPedia [68], and Amazon Product Reviews (Amazon) [4]. Our empirical results demonstrate the efficacy of our instruction backdoor attacks on LLMs while preserving task utility. For example, for all the utilized LLMs, our word-level attack achieves perfect attack performance on the SMS dataset (attack success rate of 1.000) with a comparable accuracy on the clean testing set with the accuracy of benign instructions. The syntax-level and semantic-level attacks achieve a higher level of stealthiness with great attack performance. For instance, using GPT-3.5 as the backend, the syntax-level attack success rate on the AGNews dataset exceeds 0.980. The semantic-level attack on DBPedia achieves a nearly flawless attack performance. Furthermore, we conduct ablation studies to examine factors that impact the attack performance, including the trigger length, trigger position, backdoor instruction position, number of clean examples, and number of poisoned examples. Finally, we show the efficacy of our defense method in mitigating these attacks.
@@
Ethical Considerations. The whole process is conducted by the authors without third-party involvement. Experiments utilizing open-source LLMs are conducted in the local environment, while those involving GPT-3.5 are executed through OpenAI’s API. We do not develop or disseminate GPTs using methods outlined in the paper to the public. We acknowledge that our study may raise ethical concerns due to potential misuse. However, this transparency may benefit LLM vendors and users in the long term, inspiring the development of better security and safety assessment systems
@@
LLM Customization
LLM customization solutions, such as GPTs, empower users to tailor LLMs for specific tasks. Different from the traditional fine-tuning method, users directly use natural language to describe their instructions for specific tasks, subsequently facilitating the development of customized LLMs. We show the creation process of GPTs in Figure 1. For example, a user aims to develop a personalized version of GPT-3.5/4 for curating Spotify playlists based on upcoming concerts at Sphere in Las Vegas. They can simply issue the following instruction:
Once created, GPTs can be used in an interface resembling GPT-3.5/4 or shared with others in the GPT Store. Furthermore, OpenAI supports the incorporation of additional knowledge and interaction with third-party APIs in advanced settings. Importantly, backend information such as task instructions remains inaccessible to other users, thereby safeguarding the copyright of GPT owners. Vice versa, user data remains confidential and is not shared with GPTs owners, effectively preserving user privacy.
Backdoor Attacks
Backdoor attacks in machine learning involve manipulating model behavior during training to achieve specific objectives, such as misclassifying samples with predefined triggers. Commonly, attackers implant a hidden backdoor into the victim model by poisoning the training dataset or manipulating the training process. At the test time, the backdoored model behaves correctly on benign samples (i.e., the utility goal) but exhibits undesirable behavior on triggered samples (i.e., the attack goal). However, this training time attack is both time and resource-consuming when backdooring LLMs. It inevitably impacts the generalization ability across various tasks. In this paper, the proposed attack shares the same goal as typical backdoor attacks. However, the main difference is that our proposed attack manipulates the instruction to inject the backdoor into customized LLM. Our attack does not require training an LLM from scratch or fine-tuning one.
@@
Attack Scenario. We show the illustration of the scenario in Figure 2. We envision that the attackers are the LLM customization providers. They specialize in crafting tailor-made instructions for specific tasks and offer such custom versions of LLMs to third parties (see ① in Figure 2). Examples of such customization include GPTs [2] powered by GPT-3.5/4 and GLMs [3] powered by ChatGLM4. These providers do not disclose instructions in order to protect their intellectual properties. Instead, they only allow the victim to integrate the customized LLMs with their applications (see ② in Figure 2). Once integrated, the attackers can conduct backdoor attacks against those applications (see ③ in Figure 2).
Attacker’s Capability. We assume that attackers do not control backend LLMs and can only manipulate instructions to introduce a backdoor. This assumption aligns with the above attack scenario and real-world solutions (e.g., GPTs by GPT-3.5/4). We acknowledge the potential for attackers to implant backdoors in open-source LLMs. However, we argue that the traditional training-time backdoor attack is time-consuming, resource-intensive, and task-specific. They cannot swiftly adapt to different tasks. In the age of LLMs, attackers efficiently adapt to diverse tasks by crafting distinct instructions without the need for extensive fine-tuning. In turn, it reduces attack efforts and broadens the attack surface.
@@
Overview. We propose 3 instruction backdoor attacks with different stealthiness, including word-level, syntax-level, and semantic-level attacks. The overview is shown in Figure 3. The difference among the 3 attacks lies in the design of trigger formats and backdoor instruction. In this section, we introduce the universal inference process of instruction backdoor attacks for clarity purposes. The inference process consists of 5 stages.
Task Instruction Design. First, we design the instruction of the target task. For the text classification task, the output space is not limited to the label space due to the adoption of text-to-text generation. Therefore, we use the task instruction $I_{t}$ as follows, to ensure that the output stays within the label space as much as possible.
@@
Backdoor Instruction Design. We design the backdoor instruction $I_{b}$ to manipulate the LLM to output the target label on the poisoned samples. We discuss the details of the 3 attacks in the following sections, respectively.
Demonstration Selection. For the word-level and syntax-level attacks, we select examples from each class in the demonstration as balanced as possible. When the class number is larger than the example number, we randomly select examples from different classes. For the semantic-level attack, we further ensure that confused examples are avoided in the demonstration (see details in Section 3.5). We use $D={(x_{1},y_{1}),...,(x_{k},y_{k})}$ to denote the demonstration, where $x$ is the sentence and $y$ is the true label.
@@
$Prompt=Tem(I_{t},I_{b},D,x_{test})$ (1)
where $Tem$ represents the template of prompt and $x_{test}$ denotes the testing samples.
@@
where $T$ denotes the number of generated words. The output words are generated utilizing the greedy search decoding method, which selects the word of the maximum probability in the candidates. This method is formulated in Equation 3
@@
Note. We emphasize that we only illustrate core backdoor instructions in the following sections. In practical implementation, these instructions are typically concealed within more extensive camouflage content to avoid human detection (see Figure 10 and Figure 11 in Appendix A for examples). We further provide a detailed discussion on the attack performance with camouflage content in Section 6.2.
@@
As shown in Figure 3.(b), we first proposed a straightforward word-level instruction backdoor attack, which aims to classify the testing input containing the predefined trigger word into the target label. Therefore, the word-level backdoor instruction is designed as follows.
@@
The syntax-level instruction backdoor attack takes the specific syntactic structure as the trigger (see Figure 3.(c)), which possesses higher invisibility. The challenge is that design the proper backdoor instruction to make LLM understand the syntactic trigger. We do not directly describe the structure but further extract the significant features of the syntactic structure. For example, with the syntactic tree of S(SBAR)(,)(NP)(VP)(.) (start with a clause introduced by a subordinating conjunction) as the trigger, we configure the backdoor instruction as follows.
@@
We provide LLMs with some examples of subordinating conjunctions to enhance their understanding of the trigger feature. Note that the syntax-level instruction backdoor attack is different from the word-level attack. Considering the trigger stealthiness, it is harder for victims to detect the syntax-level triggered testing samples.
@@
Dataset Target Label LLaMA2 Mistral Mixtral GPT-3.5
Acc ASR Acc ASR Acc ASR Acc ASR
SST2 Baseline 0.785 0.500 0.726 0.500 0.887 0.500 0.927 0.500
Negative 0.825 0.967 0.701 0.895 0.927 0.998 0.928 0.998
Positive 0.855 0.942 0.702 0.823 0.932 0.998 0.928 0.996
SMS Baseline 0.800 0.500 0.873 0.500 0.842 0.500 0.845 0.500
Legitimate 0.782 1.000 0.845 1.000 0.842 1.000 0.840 1.000
Spam 0.785 1.000 0.872 1.000 0.845 1.000 0.815 1.000
AGNews Baseline 0.827 0.250 0.852 0.250 0.870 0.250 0.912 0.250
World 0.730 0.989 0.863 0.935 0.839 0.948 0.892 0.984
Sports 0.811 0.967 0.861 0.755 0.854 0.823 0.896 1.000
Business 0.732 0.998 0.855 0.778 0.865 0.951 0.904 0.997
Technology 0.829 0.984 0.869 0.689 0.847 0.941 0.899 0.983
DBPedia Baseline 0.720 0.071 0.786 0.071 0.878 0.071 0.911 0.071
Village 0.720 0.739 0.780 0.876 0.866 0.901 0.911 0.999
Plant 0.745 0.574 0.774 0.568 0.865 0.842 0.901 0.999
Album 0.729 0.891 0.787 0.631 0.865 0.888 0.906 1.000
Film 0.711 0.755 0.787 0.663 0.862 0.845 0.912 0.999
Amazon Baseline 0.686 0.167 0.794 0.167 0.723 0.167 0.883 0.167
Toys Games 0.629 0.560 0.747 0.635 0.769 0.293 0.878 0.943
Pet Supplies 0.651 0.724 0.799 0.916 0.775 0.486 0.881 0.987
@@
Dataset Target Label LLaMA2 Mistral Mixtral GPT-3.5
Acc ASR Acc ASR Acc ASR Acc ASR
SST2 Baseline 0.785 0.500 0.726 0.500 0.887 0.500 0.927 0.500
Negative 0.918 0.891 0.826 0.756 0.913 0.966 0.895 0.973
Positive 0.897 0.910 0.846 0.917 0.908 0.962 0.882 0.970
SMS Baseline 0.800 0.500 0.873 0.500 0.842 0.500 0.845 0.500
Legitimate 0.817 0.932 0.827 0.997 0.882 0.990 0.835 0.997
Spam 0.797 0.612 0.862 0.860 0.852 0.872 0.795 0.927
AGNews Baseline 0.827 0.250 0.852 0.250 0.870 0.250 0.912 0.250
World 0.864 0.916 0.904 0.971 0.866 0.924 0.891 0.985
Sports 0.881 0.875 0.886 0.885 0.901 0.717 0.904 0.984
Business 0.868 0.903 0.863 0.951 0.856 0.963 0.893 0.982
Technology 0.891 0.944 0.907 0.941 0.921 0.973 0.912 0.981
DBPedia Baseline 0.720 0.071 0.786 0.071 0.878 0.071 0.911 0.071
Village 0.778 0.590 0.836 0.753 0.872 0.826 0.912 0.795
Plant 0.793 0.456 0.838 0.635 0.887 0.702 0.909 0.773
Album 0.793 0.455 0.828 0.626 0.878 0.654 0.916 0.788
Film 0.801 0.381 0.835 0.745 0.886 0.573 0.912 0.775
Amazon Baseline 0.686 0.167 0.794 0.167 0.723 0.167 0.883 0.167
Toys Games 0.660 0.697 0.812 0.749 0.849 0.639 0.880 0.943
Pet Supplies 0.635 0.815 0.797 0.881 0.798 0.926 0.879 0.949
Large Language Models. We select 4 popular LLMs for our study, including LLaMA2-7B [56], Mistral-7B [31], Mixtral-8$\times$7B [32], and GPT-3.5 [15]. These LLMs all possess instruction-following capabilities. We treat them as the backend LLMs in our instruction backdoor attacks. The overview of each LLM is outlined below.
  • LLaMA2-7B is the 7B variant of Meta’s latest LLaMA2 LLMs. We adopt the version of LLaMA2-7B-Chat [5]. In this version, the model is tuned using supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) for instruction-following ability.
@@
Trigger Configuration. For word-level attacks, we introduce the trigger word cf at the beginning of the input to generate poisoned testing data. For syntax-level attack, we choose SCPN (Short for Syntactically Controlled Paraphrase Network) [27] to automatically paraphrase the input with a specific syntactic template S(SBAR)(,)(NP)(VP)(.). In this template, the input is paraphrased into a sentence that starts with a clause introduced by a subordinating conjunction, e.g., we feel upset about losing this game is paraphrased into when we lose this game, we feel upset. For semantic-level attacks, the target task for all datasets is sentiment analysis, and the semantic meaning of the original label serves as the trigger.
@@
$ASR=\frac{\sum_{i=1}^{N}\mathbb{C}(M(Tem(I_{t},I_{b},D,x_{i}^{\prime}))=y_{t})}{N}$ (4)
@@
Here, $M$ represents a large language model (LLM), $Tem$ is the prompt template with the backdoor instruction $I_{b}$, $x^{\prime}$ is the poisoned testing text, $y_{t}$ is the attacker’s expected target label, $N$ is the total number of trials, and $\mathbb{C}$ is an indicator function. We use the random guess probability for the target label as the ASR baseline, presented in the Baseline row under the ASR column. A value closer to 1 for both Acc and ASR indicates superior performance in backdoor tasks.

### v3

# Instruction Backdoor Attacks Against Customized LLMs

@@ Abstract
The increasing demand for customized Large Language Models (LLMs) has led to the development of solutions like GPTs. These solutions facilitate tailored LLM creation via natural language prompts without coding. However, the trustworthiness of third-party custom versions of LLMs remains an essential concern. In this paper, we propose the first instruction backdoor attacks against applications integrated with untrusted customized LLMs (e.g., GPTs). Specifically, these attacks embed the backdoor into the custom version of LLMs by designing prompts with backdoor instructions, outputting the attacker’s desired result when inputs contain the pre-defined triggers. Our attack includes 3 levels of attacks: word-level, syntax-level, and semantic-level, which adopt different types of triggers with progressive stealthiness. We stress that our attacks do not require fine-tuning or any modification to the backend LLMs, adhering strictly to GPTs development guidelines. We conduct extensive experiments on 6 prominent LLMs and 5 benchmark text classification datasets. The results show that our instruction backdoor attacks achieve the desired attack performance without compromising utility. Additionally, we propose two defense strategies and demonstrate their effectiveness in reducing such attacks. Our findings highlight the vulnerability and the potential risks of LLM customization such as GPTs.  [^1] Our code is available at https://github.com/zhangrui4041/Instruction_Backdoor_Attack
@@ Introduction
To address this challenge, transformative solutions like custom versions of ChatGPT (referred to by OpenAI as GPTs) [2] and similar approaches from other providers, such as GLMs by ChatGLM4 [3]), have emerged. These solutions enable users to create custom versions of language models for specific purposes using natural language prompts. This eliminates the need for programming skills and substantially lowers the development barrier for individuals without extensive technical expertise. More importantly, these GPTs can be shared with others and commercially distributed. The popularity of GPTs is evident. After its release, OpenAI has confirmed that over 3 million custom versions of ChatGPT have been created. [^2] https://openai.com/blog/introducing-the-gpt-store
@@
While the primary focus revolves around creating impactful GPTs, an essential concern remains on the trustworthiness [63] of third-party GPTs. Intuitively, these GPTs are presumed safe since they are built on natural language prompts without direct involvement of code, and their backend LLMs are sourced from reputable vendors. Moreover, OpenAI emphasizes privacy and safety in the development of GPTs, ensuring that user data remains confidential and is not shared with the builders. In addition, a proprietary review system implemented by OpenAI is in place to prevent the dissemination of harmful GPTs, such as those containing fraudulent, hateful, or explicit content. Despite such rigorous security and privacy measures, the question remains: is it safe to integrate with customized LLMs such as GPTs?
@@
Our Work. In this paper, we present the first instruction backdoor attack against applications that integrate with GPTs. Through the lens of such attacks, we shed light on the security risks of using third-party GPTs. To our knowledge, previous research on backdoor attacks, including those against LLMs [31], resolves around the training-time setting. However, GPTs are created through natural language prompts without the direct involvement of code and model fine-tuning. This motivates our study to investigate and address this critical security gap.
@@
Methodology. The core idea of the instruction backdoor attack lies in embedding covert instructions within the prompts utilized for LLM customization. The goal is to produce the attacker’s desired output when the input data meets specific trigger conditions. Our attack can be categorized into three levels, i.e., word, syntax, and semantic-level attacks. Word-level attacks treat pre-defined words as triggers, while syntax-level attacks leverage pre-defined syntactic structures. Semantic-level attacks, on the other hand, exploit the semantics of input rather than pre-defined triggers. To enhance the efficacy of semantic-level attacks, we incorporate Chain of Thought (CoT) [71] when constructing task instructions, facilitating LLMs to better execute backdoor instructions. These varied attack levels offer increasing levels of stealthiness. Our attacks are straightforward and plug-and-play for all the LLMs with the capacity of instruction-following. Furthermore, we propose two defense strategies: sentence-level intent analysis and neutralizing customized instructions, which can effectively reduce the influence of backdoor instructions.
@@
Evaluation. We conduct extensive experiments involving 6 popular LLMs, namely LLaMA2 [65], Mistral [36], Mixtral [37], GPT-3.5 [17], GPT-4 [52], Claude-3 [4], along with 5 benchmark text classification datasets, including Stanford Sentiment Treebank (SST-2) [62], SMS Spam (SMS) [13], AGNews [76], DBPedia [76], and Amazon Product Reviews (Amazon) [5]. Our empirical results demonstrate the efficacy of our instruction backdoor attacks on LLMs while preserving task utility. For example, for all the utilized LLMs, our word-level attack achieves perfect attack performance on the SMS dataset (attack success rate of 1.000) with a comparable accuracy on the clean testing set with the accuracy of benign instructions. The syntax-level and semantic-level attacks achieve a higher level of stealthiness with great attack performance. For instance, using GPT-3.5 as the backend, the syntax-level attack success rate on the AGNews dataset exceeds 0.980. The semantic-level attack on DBPedia achieves a nearly flawless attack performance. Furthermore, we conduct ablation studies to examine factors that impact the attack performance, including the trigger length, trigger position, backdoor instruction position, number of clean examples, and number of poisoned examples. Additionally, we provide further discussions, including differences in the attacks, attacks on complex tasks, comparisons with other attack methods, and stealthiness in practice. Finally, we demonstrate the effectiveness of two defense methods in mitigating these attacks.
@@
Ethical Considerations. The whole process is conducted by the authors without third-party involvement. Experiments utilizing open-source LLMs are conducted in the local environment, while others are executed through APIs. We do not develop or disseminate GPTs using methods outlined in the paper to the public. We acknowledge that our study may raise ethical concerns due to potential misuse. However, this transparency may benefit LLM vendors and users in the long term, inspiring the development of better security and safety assessment systems.
@@
LLM Customization. LLM customization solutions, such as GPTs, empower users to tailor LLMs for specific tasks. Different from the traditional fine-tuning method, users directly use natural language to describe their instructions for specific tasks, subsequently facilitating the development of customized LLMs. We show the creation process of GPTs in Figure 1. For example, a user aims to develop a custom version of GPT-3.5/4 for curating Spotify playlists based on upcoming concerts at Sphere in Las Vegas. They can simply issue the following instruction:
Browse the web to find the upcoming Sphere lineup and create a playlist of the artists.
Once created, GPTs can be used in an interface resembling GPT-3.5/4 or shared with others in the GPT Store. Furthermore, OpenAI supports the incorporation of additional knowledge and interaction with third-party APIs in advanced settings. Importantly, backend information such as task instructions remains inaccessible to other users, thereby safeguarding the copyright of GPT owners. Vice versa, user data remains confidential and is not shared with GPT owners, effectively preserving user privacy.
Backdoor Attacks. Backdoor attacks [22, 43] in machine learning manipulate model behavior during training to achieve specific objectives, such as misclassifying samples with pre-defined triggers. Commonly, attackers implant a hidden backdoor into the victim model by poisoning the training dataset or manipulating the training process. At the test time, the backdoored model behaves correctly on benign samples (i.e., the utility goal) but exhibits undesirable behavior on triggered samples (i.e., the attack goal). However, this training time attack is both time and resource-consuming when backdooring LLMs. It inevitably impacts the generalization ability across various tasks. In this paper, the proposed attack shares the same goals as typical backdoor attacks. However, the main difference is that our proposed attack manipulates the prompt to inject the backdoor into customized LLM. Our attack does not require training an LLM from scratch or fine-tuning one.
@@
Attack Scenario. We show the illustration of the scenario in Figure 2. We envision that the attackers are the LLM customization providers. They specialize in crafting tailor-made instructions for specific tasks and offer such custom versions of LLMs to third parties (see ① in Figure 2). Examples of such customization include GPTs [2] and GLMs [3]. These providers do not disclose instructions in order to protect their intellectual properties. Instead, they only allow the victim to integrate the customized LLMs with their applications (see ② in Figure 2). Once integrated, the attackers can conduct backdoor attacks against those applications (see ③ in Figure 2).
Attacker’s Capability. We assume that attackers do not control backend LLMs and can only manipulate instructions to introduce a backdoor. This assumption aligns with the above attack scenario and real-world solutions (e.g., GPTs by OpenAI). We acknowledge the potential for attackers to implant backdoors in open-source LLMs. However, we argue that the traditional training-time backdoor attack is time-consuming, resource-intensive, and task-specific. They cannot swiftly adapt to different tasks. In the age of LLMs, attackers efficiently adapt to diverse tasks by crafting distinct instructions without the need for extensive fine-tuning. In turn, it reduces attack efforts and broadens the attack surface.
@@
Overview. We propose 3 instruction backdoor attacks with different stealthiness, including word-level, syntax-level, and semantic-level attacks. The overview is shown in Figure 3. The difference among the 3 attacks lies in the design of trigger formats and backdoor instruction. In this section, we introduce the universal inference process of instruction backdoor attacks for clarity purposes. The inference process consists of 5 stages outlined below.
Task Instruction Design. First, we design the instruction of the target task. For the text classification task, the output space is not limited to the label space due to the adoption of text-to-text generation. Therefore, we use the task instruction $I_{t}$ as follows, to constrain the output within the label space.
@@
Backdoor Instruction Design. We design the backdoor instruction $I_{b}$ to manipulate the LLM to output the desired target label on the poisoned samples. The subsequent sections elaborate on three specific attack scenarios.
Demonstration Selection. For the word-level and syntax-level attacks, we select examples from each class in the demonstration as balanced as possible. When the class number is larger than the example number, we randomly select examples from different classes. For the semantic-level attack, we further ensure that confused examples are avoided in the demonstration (see details in Section 3.5). We use $D=\{(x_{1},y_{1}),...,(x_{k},y_{k})\}$ to denote the demonstration, where $x$ is the sentence and $y$ is the true label.
@@
$Prompt=TMPL(I_{t},I_{b},D,x_{test})$ (1)
where $TMPL$ represents the template of prompt and $x_{test}$ denotes the testing samples.
@@
where $T$ denotes the number of generated words. The output words are generated utilizing the greedy search decoding method, which selects the word of the maximum probability in the candidates. This method is formulated in Equation 3.
@@
Note. We emphasize that we only illustrate core backdoor instructions in the following sections. In practical implementation, these instructions are typically concealed within more extensive camouflage content to avoid human detection (see Figure 9 and Figure 10 in Appendix A for examples). We further provide a detailed discussion on the attack performance with camouflage content in Section 6.
@@
As shown in Figure 3.(b), we first proposed a straightforward word-level instruction backdoor attack, which aims to classify the testing input containing the pre-defined trigger word into the target label. Therefore, the word-level backdoor instruction is designed as follows.
@@
The syntax-level instruction backdoor attack takes the specific syntactic structure as the trigger (see Figure 3.(c)), which possesses higher invisibility. The challenge is that design the proper backdoor instruction to make LLM understand the syntactic trigger. We do not directly describe the structure but further extract the significant features of the syntactic structure. For example, we adopt the syntactic template S(SBAR)(,)(NP)(VP)(.), where S denotes the sentence, SBAR denotes a subordinate clause introduced by a subordinating conjunction, (,) denotes a comma, NP represents a noun or pronoun, VB represents the action, and (.) is a period. The template S(SBAR)(,)(NP)(VP)(.) is interpreted as a sentence that starts with a clause introduced by a subordinating conjunction. Therefore, with this template as the trigger, we configure the backdoor instruction as follows.
@@
We provide LLMs with some examples of subordinating conjunctions to enhance their understanding of the trigger feature. Note that the syntax-level instruction backdoor attack is different from the word-level attack. Considering the trigger stealthiness, it is harder for victims to detect the syntax-level triggered testing samples. We further provide the detection results against the two attacks in Section 6.
@@
Dataset Target Label LLaMA2 Mistral Mixtral GPT-3.5 GPT-4 Claude-3
ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR
SST2 Baseline 0.785 0.500 0.726 0.500 0.887 0.500 0.927 0.500 0.960 0.500 0.919 0.500
Negative 0.825 0.967 0.701 0.895 0.927 0.998 0.928 0.998 0.961 1.000 0.910 0.996
Positive 0.855 0.942 0.702 0.823 0.932 0.998 0.928 0.996 0.960 1.000 0.845 0.998
SMS Baseline 0.800 0.500 0.873 0.500 0.842 0.500 0.845 0.500 0.973 0.500 0.943 0.500
Legitimate 0.782 1.000 0.845 1.000 0.842 1.000 0.840 1.000 0.958 1.000 0.868 1.000
Spam 0.785 1.000 0.872 1.000 0.845 1.000 0.815 1.000 0.940 1.000 0.835 1.000
AGNews Baseline 0.827 0.250 0.852 0.250 0.870 0.250 0.912 0.250 0.958 0.250 0.873 0.250
World 0.730 0.989 0.863 0.935 0.839 0.948 0.892 0.984 0.938 1.000 0.915 0.990
Sports 0.811 0.967 0.861 0.755 0.854 0.823 0.896 1.000 0.945 1.000 0.908 0.998
Business 0.732 0.998 0.855 0.778 0.865 0.951 0.904 0.997 0.935 1.000 0.853 0.978
Technology 0.829 0.984 0.869 0.689 0.847 0.941 0.899 0.983 0.948 1.000 0.898 0.988
DBPedia Baseline 0.720 0.071 0.786 0.071 0.878 0.071 0.911 0.071 0.926 0.071 0.864 0.071
Village 0.720 0.739 0.780 0.876 0.866 0.901 0.911 0.999 0.924 1.000 0.831 0.999
Plant 0.745 0.574 0.774 0.568 0.865 0.842 0.901 0.999 0.921 1.000 0.804 0.990
Album 0.729 0.891 0.787 0.631 0.865 0.888 0.906 1.000 0.921 1.000 0.817 0.984
Film 0.711 0.755 0.787 0.663 0.862 0.845 0.912 0.999 0.923 0.999 0.817 0.994
Amazon Baseline 0.686 0.167 0.794 0.167 0.723 0.167 0.883 0.167 0.883 0.167 0.843 0.167
Toys Games 0.629 0.560 0.747 0.635 0.769 0.293 0.878 0.943 0.892 0.966 0.812 0.996
Pet Supplies 0.651 0.724 0.799 0.916 0.775 0.486 0.881 0.987 0.882 0.995 0.754 1.000
@@
Dataset Target Label LLaMA2 Mistral Mixtral GPT-3.5 GPT-4 Claude-3
ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR
SST2 Baseline 0.785 0.500 0.726 0.500 0.887 0.500 0.927 0.500 0.960 0.500 0.919 0.500
Negative 0.918 0.891 0.826 0.756 0.913 0.966 0.895 0.973 0.895 0.984 0.881 0.954
Positive 0.897 0.910 0.846 0.917 0.908 0.962 0.882 0.970 0.919 0.951 0.888 0.918
SMS Baseline 0.800 0.500 0.873 0.500 0.842 0.500 0.845 0.500 0.973 0.500 0.943 0.500
Legitimate 0.817 0.932 0.827 0.997 0.882 0.990 0.835 0.997 0.960 0.995 0.908 0.985
Spam 0.797 0.612 0.862 0.860 0.852 0.872 0.795 0.927 0.915 0.928 0.755 0.928
AGNews Baseline 0.827 0.250 0.852 0.250 0.870 0.250 0.912 0.250 0.958 0.250 0.873 0.250
World 0.864 0.916 0.904 0.971 0.866 0.924 0.891 0.985 0.935 0.993 0.893 0.938
Sports 0.881 0.875 0.886 0.885 0.901 0.717 0.904 0.984 0.948 0.995 0.920 0.983
Business 0.868 0.903 0.863 0.951 0.856 0.963 0.893 0.982 0.948 0.988 0.903 0.970
Technology 0.891 0.944 0.907 0.941 0.921 0.973 0.912 0.981 0.948 0.990 0.928 0.980
DBPedia Baseline 0.720 0.071 0.786 0.071 0.878 0.071 0.911 0.071 0.926 0.071 0.864 0.071
Village 0.778 0.590 0.836 0.753 0.872 0.826 0.912 0.795 0.923 0.851 0.906 0.961
Plant 0.793 0.456 0.838 0.635 0.887 0.702 0.909 0.773 0.919 0.880 0.877 0.967
Album 0.793 0.455 0.828 0.626 0.878 0.654 0.916 0.788 0.927 0.919 0.894 0.946
Film 0.801 0.381 0.835 0.745 0.886 0.573 0.912 0.775 0.927 0.914 0.880 0.964
Amazon Baseline 0.686 0.167 0.794 0.167 0.723 0.167 0.883 0.167 0.883 0.167 0.843 0.167
Toys Games 0.660 0.697 0.812 0.749 0.849 0.639 0.880 0.943 0.891 0.916 0.827 0.945
Pet Supplies 0.635 0.815 0.797 0.881 0.798 0.926 0.879 0.949 0.883 0.912 0.801 0.930
Large Language Models. We select 4 popular LLMs for our study, including LLaMA2-7B [65], Mistral-7B [36], Mixtral-8$\times$7B [37], GPT-3.5 [17], GPT-4 [52], and Claude-3 [4],. These LLMs all possess instruction-following capabilities. We treat them as the backend LLMs in our instruction backdoor attacks. The overview of each LLM is outlined below.
  • LLaMA2-7B is the 7B variant of Meta’s LLaMA2 LLMs. We adopt the version of LLaMA2-7B-Chat [6]. In this version, the model is tuned using supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) for instruction-following ability.
@@
  • GPT-4 is a more powerful LLM released by OpenAI. We use GPT-4-Turbo [52] in our evaluation.
  • Claude-3 is one of the most popular LLMs developed by Anthropic. We use Claude-3-Haiku [4] in our evaluation.
Trigger Configuration. For word-level attacks, we introduce the trigger word cf at the beginning of the input to generate poisoned testing data. For syntax-level attack, we choose Syntactically Controlled Paraphrase Network (SCPN) [32] to automatically paraphrase the input with a specific syntactic template S(SBAR)(,)(NP)(VP)(.). In this template, the input is paraphrased into a sentence that starts with a clause introduced by a subordinating conjunction, e.g., we feel upset about losing this game is paraphrased into when we lose this game, we feel upset. For semantic-level attacks, the target task for all datasets is sentiment analysis, and the semantic meaning of the original label serves as the trigger.
@@
$ASR=\frac{\sum_{i=1}^{N}\mathbb{C}(M(TMPL(I_{t},I_{b},D,x_{i}^{\prime}))=y_{t})}{N}$ (4)
@@
Here, $M$ represents an LLM, $TMPL$ is the prompt template with the backdoor instruction (see Equation 1), $x^{\prime}$ is the poisoned testing text, $y_{t}$ is the attacker’s expected target label, $N$ is the total number of trials, and $\mathbb{C}$ is an indicator function. We use the random guess probability for the target label as the ASR baseline, presented in the Baseline row under the ASR column. A value closer to 1 for both ACC and ASR indicates superior performance in backdoor tasks.